# Simulating Longitudinal Data from Marginal Structural Models

### v1

@@ 3.2.2 Simulating from pair-copula-constructions
Proof.
[…]
$=\frac{\partial}{\partial F(x_{j}\,|\,\bm{x}_{-j})}C_{YX_{j}\mid\bm{x}_{-j}}\{F(y\,|\,\bm{x}_{-j}),F(x_{j}\,|\,\bm{x}_{-j})\}.$
For convenience of notation, we introduce the h-function $h(x_{1},x_{2};\Theta)$ from Aas et al. (2009). Let $x_{1}$ and $x_{2}$ be copula random variables, i.e. with margins following the uniform distribution $x_{i}\sim\operatorname{U}(0,1)$ and $F_{i}(x_{i})=x_{i}$, for $i=1,2$. The $h$-function is defined as:
$h(x_{1},x_{2};\Theta)$ $:=F(x_{1}\,|\,x_{2})=\frac{\partial F_{12}(x_{1},x_{2})}{\partial x_{2}f_{2}(x_{2})}$
$=\frac{\partial}{\partial F_{2}(x_{2})}C_{12}(F_{1}(x_{1}),F_{2}(x_{2});\Theta)$ (6)
$=\frac{\partial}{\partial x_{2}}C_{12}(x_{1},x_{2};\Theta),$ (7)
where the second argument in $h(\cdot;\Theta)$ is the conditional variable, and $\Theta$ represents the set of copula parameters. From (7) we can see that the $h$-function is essentially the conditional copula function, $C(u_{1}\,|\,u_{2};\Theta)$.
Example 3.1.
The derivation of the analytical form of $h(\cdot)$ for common bivariate copulas is straightforward following (7) for parametric copula families. Some examples include:
  • Gaussian copula: $h(x_{1},x_{2};\rho)=\Phi\left(\frac{\displaystyle\Phi^{-1}(x_{1})-\rho\Phi^{-1}(x_{2})}{\displaystyle\sqrt{1-\rho^{2}}}\right)$;
  • Student’s $t$ copula: $h(x_{1},x_{2};\rho,\nu)=t_{\nu+1}\left(\frac{\displaystyle\sqrt{\nu+1}(t_{\nu}^{-1}(x_{1})-\rho t_{\nu}^{-1}(x_{2}))}{\displaystyle\sqrt{\left(\nu+\left(t_{\nu}^{-1}(x_{2})\right)^{2}\right)\left(1-\rho^{2}\right)}}\right)$;
  • Clayton copula: $h(x_{1},x_{2};\theta)=x_{2}^{-\theta-1}\left(x_{1}^{-\theta}+x_{2}^{-\theta}-1\right)^{-1-1/\theta}$;
  • Gumbel copula: $h(x_{1},x_{2};\theta)=C(x_{1},x_{2};\theta)\frac{1}{x_{2}}(-\log x_{2})^{\theta-1}\left((-\log x_{1})^{\theta}+(-\log x_{2})^{\theta}\right)^{1/\theta-1}$.
The $h$-functions are foundational for simulating from $F(y\,|\,\bm{x})$. Similar to (4), by imposing an order $1,\dots,d$ on the components of $\bm{x}$, we can recursively apply $h$-functions to obtain:
$F(y\,|\,\bm{x})$ $=h\{F(y\,|\,\bm{x}_{1:{d-1}}),F(x_{d}\,|\,\bm{x}_{1:{d-1}}),\Theta_{1}\}$
$=h[h\{F(y\,|\,\bm{x}_{1:{d-2}}),F(x_{d-1}\,|\,\bm{x}_{1:{d-2}};\Theta_{2}\},F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{1}]$
$=h\left[h\{h\left(F(y\,|\,\bm{x}_{1:{d-3}}),F(x_{d-2}\,|\,\bm{x}_{1:{d-3}};\Theta_{3})\right),F(x_{d-1}\,|\,\bm{x}_{1:{d-2})};\Theta_{2}\},F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{1}\right]$
$\dots$
$=h[\dots h\{h(F(y),F(x_{1});\Theta_{d}),F(x_{2}\,|\,x_{1});\Theta_{d-1}\}\dots;\Theta_{1}].$ (8)
Thus, $Y$ given $\bm{X}$ can be simulated by first drawing a uniform $U\sim{\rm U}(0,1)$, and then evaluating:
$Y=h^{-1}[\dots h^{-1}\{h^{-1}(U,\,F(X_{1});\Theta_{d}),F(X_{2}\,|\,X_{1});\Theta_{d}\}\dots;\Theta_{1}].$ (9)
This means that to simulate $Y$ given a $d$-dimensional vector $\bm{X}$, we recursively calculate the $d$ inverse $h$-functions, forming a PCC. This forms the basis of our proposed simulation method.
@@ 4 Simulation Algorithm
Figure 1: An example directed acyclic graph showing temporal relationships in a longitudinal discrete setting, with baseline covariates $Z$; treatments $A_{0},A_{1},\ldots,A_{K}$; time-varying covariates $L_{1},\ldots,L_{K}$ and outcomes $Y$ observed at the end of the K follow up intervals. Given the topological ordering $Z,L_{0},A_{0},L_{1},A_{1}\ldots,L_{K},A_{K},Y$, any variable may depend upon any subset of the previous variables.
We consider a longitudinal study with $n$ individuals observed over discrete time points, $\mathcal{K}=\{0,1,\dots,K\}$. At each follow-up $k$, treatment status $A_{k}$, and a set of time-varying covariates $L_{k}$ are recorded. Additionally, $Z$ denotes a set of static covariates measured at baseline. There may also be latent variables, represented by $\{U_{k}\}_{k\in\mathcal{K}}$. To indicate the history of a variable up to time $k$, we use a bar notation. For instance, $\bar{A}_{k}=\{A_{0},A_{1}\dots,A_{k}\}$ and $\bar{L}_{k}=\{L_{0},L_{1},\dots,L_{k}\}$. For convenience, we let $\bar{A}_{-1}=\emptyset$ and $\bar{L}_{-1}=\emptyset$. The outcome $Y$ is observed at the final time point $K$. Figure 1 shows a DAG representing this setup, which is common in medical and epidemiological research, such as examining the effects of iron supplement use during pregnancy on the odds of anemia at delivery (Bodnar et al., 2004), anti-diabetic drugs on weight loss over a specified period (e.g. Jastreboff et al., 2022; Ryan et al., 2024) and exposure to community violence on youth’s mental health (Kennedy et al., 2023). In Section 5, we will illustrate our method with a simulation study that replicates such setups.
@@
4.1 Fully conditional approach
For simplicity, we initially assume the time-varying covariate $L_{k}$ is one-dimensional and exclude the latent process $\{U_{k}\}_{k\in\mathcal{K}}$; the extension to multi-dimensional time-varying covariates and latent variables is discussed later in this section.
The usual strategy for simulation of longitudinal samples $(Z,\bar{L}_{K},\bar{A}_{K},Y)$ is by time step in blocks using the Rosenblatt transformation (Rosenblatt, 1952). Specifically:
@@ 4.1 Fully conditional approach
    Baseline simulation: Simulate static baseline variables $z=F_{Z}^{-1}(w_{z})$, where $w_{z}$ is a uniform random variable.
@@
    Time-step simulation: For each time point $k=0,1,\dots,K$, simulate:
    $l_{k}$ $=F_{L_{k}}^{-1}(w_{k,1},\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$ (10)
    $a_{k}$ $=F_{A_{k}}^{-1}(w_{k,2},\,|\,z,\bar{l}_{k},\bar{a}_{k-1}),$
    where $w_{k,1},w_{k,2}$ are random uniform variables.
@@
    Outcome simulation: After simulating all time-varying covariates and treatment, simulate the outcome $Y$ as
    $y$ $=F_{Y}^{-1}(w_{y}\,|\,z,\bar{l}_{K},\bar{a}_{K}),$ (11)
    with $w_{y}$ uniformly distributed.
@@ 4.2 Frugal Parameterization for Outcome Simulation
Our method diverges from traditional conditional approaches by using the frugal parameterization for $Y$ to avoid intractable analytical calculations for the marginal structural model. Instead of directly assuming a conditional distribution $F_{Y}^{-1}(w_{y}\,|\,z,\bar{l}_{K},\bar{a}_{K})$, we apply interventional distributions to decompose this into interventional margins and dependency structures.
Evans and Didelez (2024) showed that the conditional distribution $F_{Y}^{-1}(w_{y}\,|\,z,\bar{l}_{k},\bar{a}_{K})$ can be frugally parameterized as:
$p(y\,|\,z,\bar{l}_{k},\bar{a}_{K})=p^{*}_{Y\mid Z\bar{A}_{K}}(y\,|\,z,do(\bar{a}_{K}))\cdot c^{*}_{Y\bar{L}_{K}\mid Z\bar{A}_{K}}(u^{*}_{y},u^{*}_{l_{0}},\dots,u^{*}_{l_{K}}\,|\,z,do(\bar{a}_{K})),$ (12)
where $u^{*}_{y}=F(y\,|\,z,do(\bar{a}_{K}))$ and $u^{*}_{l_{k}}=F(l_{k}\,|\,z,do(\bar{a}_{K}))$ for $k=0,1,\ldots,K$ are the marginal quantiles under interventions that set $\bar{A}_{K}=\bar{a}_{K}$. We mark these quantiles with an asterisk to emphasize their interventional nature. This decomposition implies that we can simulate $Y$ through the interventional copula and the interventional margin. Applying the pair-copula-construction from (4) to the expression in (12), we obtain:
$p(y\,|\,z,\bar{l}_{K},\bar{a}_{K})$ $=c^{*}_{YL_{K}\mid\bar{L}_{K-1},Z,\bar{A}_{K}}\{F(y\,|\,\bar{l}_{K-1},z,do(\bar{a}_{K})),F(l_{K}\,|\,\bar{l}_{K-1},z,do(\bar{a}_{K}));\Theta_{K}\}$
$\quad\cdot c^{*}_{YL_{K-1}\mid\bar{L}_{K-2},Z,\bar{A}_{K}}\{F(y\,|\,\bar{l}_{K-2},z,do(\bar{a}_{K})),F(l_{K-1}\,|\,\bar{l}_{K-2},z,do(\bar{a}_{K}));\Theta_{K-1}\}$
[…]
$\quad\cdot c^{*}_{YL_{0}\mid Z,\bar{A}_{K}}\{F(y\,|\,z,do(\bar{a}_{K})),F(l_{0}\,|\,z,do(\bar{a}_{K}));\Theta_{0}\}\cdot p^{*}_{Y\mid Z\bar{A}_{K}}(y\,|\,z,do(\bar{a}_{K}))$ (13)
$=\prod_{k=0}^{K}c^{*}_{Y,L_{k}\mid\bar{L}_{k-1},Z,\bar{A}_{K}}\{F(y\,|\,\bar{l}_{k-1},z,do(\bar{a}_{K})),F(l_{k}\,|\,\bar{l}_{k-1},z,do(\bar{a}_{K}));\Theta_{k}\}.$ (14)
This decomposes the conditional density into the product of $K+1$ pair-copulas and the target marginal causal density. Under intervention, graphically speaking, all arrows into nodes $\bar{A}_{k}$ are removed in the directed acyclic graph, rendering $L_{k}$ independent of subsequent interventions. Consequently, $F(l_{k}\,|\,\bar{l}_{k-1},z,do(\bar{a}_{K}))=F(l_{k}\,|\,\bar{l}_{k-1},z,\bar{a}_{k-1})$ for $k=1,2,\ldots,K$ and $F(l_{0}\,|\,z,do(\bar{a}_{K}))=F(l_{0}\,|\,z)$. Given this, following the simulation strategy from pair-copula constructions as described in Section 3.2.2, the conditional cumulative distribution function in (11) becomes
$F(y\,|\,z,\bar{l}_{K},\bar{a}_{K})$ $=h[\dots h\{h(F^{*}(y\,|\,z,\bar{a}_{K}),F(l_{0}\,|\,z);\Theta_{0}),F(l_{1}\,|\,l_{0},z,a_{0});\Theta_{1}\}\dots;\Theta_{K}].$
@@
    Calculate $u^{*}_{y}$ as
    $u^{*}_{y}$ $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},F(l_{K}\,|\,\bar{l}_{K-1},z,\bar{a}_{K});\Theta_{K}),$
    $\qquad\qquad\qquad F(l_{K-1}\,|\,\bar{l}_{K-2},z,\bar{a}_{K-1});\Theta_{K-1}\big{\}}\dots,F(l_{0}\,|\,z);\Theta_{0}\big{]}$ (15)
    $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},w_{K,1};\Theta_{K}),w_{K-1,1};\Theta_{K-1}\big{\}},\dots,w_{0,1};\Theta_{0}\big{]}.$ (16)
@@
    Obtain $y$ as:
    $y=F^{*-1}(u^{*}_{y}\,|\,z,do(\bar{a}_{K})).$
The quantiles in the inverse $h$ functions, such as $F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k},z)$, are precisely the random variables $w_{k,1}$ used in simulate them in (2). This means that we can directly substitute the $l_{k}$ quantiles into (15) without the need for computation.
@@ 4.3 Simulation Algorithm
To simulate data $(Z,\bar{L}_{K},\bar{A}_{K},Y)$ in this example, the user specifies:
  • the population distribution of the static baseline covariate $Z$: $F_{Z}$;
  • the development of the time-varying confounder $L_{k}$ as influenced by $Z$ and treatment history $\bar{A}_{k-1}$: $F_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}$;
  • treatment assignment based on the time-varying confounder $L_{k}$, past treatment and baseline covariate $Z$: $F_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1}}$;
  • the target causal marginal structural model: $F_{Y_{k+1}\mid\bar{A}_{k},Z}^{*}(\cdot\,|\,do(\bar{a}_{K}),z)$;
  • $K+1$ pair-copulas describing dependencies between outcome $Y$ and $L_{0},L_{1},\ldots,L_{K}$:
    $C_{YL_{k}\mid Z,do(\bar{A}_{k}),\bar{L}_{k-1}}\quad\text{for}\quad k=0,1,\ldots,K.$
Section 5 provides examples of specifying and interpreting these inputs through a simulation study that capture real-world complexities.
@@
     Sample $w_{k,1},w_{k,2}$ independent uniform on $[0,1]$.
@@
4.4 Extension to Include Time-Varying Latent Variables
We now extend the simulation algorithm discussed above to allow for a latent process that confounds the covariates and the final outcome $Y$. As an example, the steps to simulate longitudinal data from the model depicted in Figure 2 are:
[…]
Figure 2: Extension to the model in Figure 1 in which we also allow an unobserved process $U_{t}$ affecting only $Z$, $L_{k}$ and $Y$.
  1. 1.
    Baseline simulation: The DAG in Figure 1 imposes that the latent variable $U_{0}$ precedes $Z$ in topological order, which may indicate an initial selection bias. Thus, we first simulate the baseline value of $U$ as $u_{0}=F_{U_{0}}^{-1}(w_{0,0})$, followed by the static baseline covariate $Z$ given $U_{0}$: $z=F_{Z\mid U_{0}}^{-1}(w_{z}\,|\,u_{0})$, where $w_{0,0}$ and $w_{z}$ are uniform random numbers.
    $u_{0}$ $=F_{U_{0}}^{-1}(w_{0,0})$
    $z$ $=F_{Z\mid U_{0}}^{-1}(w_{z}\,|\,u_{0})$
    $l_{0}$ $=F_{L_{0}\mid ZU_{0}}^{-1}(w_{0,1}\,|\,z,u_{0})$
    $a_{0}$ $=F_{A_{0}\mid ZU_{0}L_{0}}^{-1}(w_{0,2}\,|\,z,u_{0},l_{0})$
  2. 2.
    Time-step simulation: For each time point $k=1,\dots,K$, simulate:
    $u_{k}$ $=F_{U_{k}\mid\bar{U}_{k-1}}^{-1}(w_{k,1}\,|\,\bar{u}_{k-1})$
    $l_{k}$ $=F_{L_{k}\mid Z\bar{U}_{k}\bar{L}_{k-1}\bar{A}_{k-1}}^{-1}(w_{k,2},\,|\,z,\bar{u}_{k},\bar{l}_{k-1},\bar{a}_{k-1})$
    $a_{k}$ $=F_{A_{k}\mid Z\bar{U}_{k}\bar{L}_{k}\bar{A}_{k-1}}^{-1}(w_{k,3},\,|\,z,\bar{u}_{k},\bar{l}_{k},\bar{a}_{k-1})$
    where $w_{k,1},w_{k,2},w_{k,3}$ are random uniform variables.
  3. 3.
    Outcome simulation: After simulating all time-varying covariates and treatment, and applying the same derivation following the frugal parameterization as in Section 4.2, simulate the outcome $Y$ in two steps:
    1. (a)
      Calculate
      $u^{*}_{y}$ $=h^{-1}\big{[}\dots h^{-1}[h^{-1}\left\{h^{-1}\left(w_{y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{u}_{K},z,\bar{a}_{K});\Theta_{K,1}\right),F(u_{K}\,|\,\bar{l}_{K-1},\bar{u}_{K-1},z,\bar{a}_{K});\Theta_{K,2}\right\}$
      $F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{u}_{K-1},z,\bar{a}_{K-1});\Theta_{K-1,1}]\dots;\Theta_{0,2}\big{]}.$
    2. (b)
      Obtain: $y=F^{*^{-1}}(u^{*}_{y}\,|\,z,do(\bar{a}_{K}))$.
    Here the $w_{y}$ are independent and uniformly distributed on $(0,1)$.
In addition to the information required in the case without latent confounders, the user also specifies:
  • the development of the latent confounder $U_{k}$ as influenced by $Z$ and treatment history $\bar{A}_{k}$;
  • $K+1$ pair-copula densities describing dependencies between outcome $Y$ and each of $U_{0},U_{1},\ldots,U_{K}$:
    $c_{YU_{k}\mid Z,\bar{U}_{k-1},\bar{L}_{k-1},do(\bar{A}_{k-1})},$
    for $k=0,1,\ldots,K$.
Extending to include additional time-varying covariates follows a similar approach. We specify the evolution of each time-varying covariate and use pair copulas to parameterize their association with the outcome, conditioned on all preceding variables. With common choices of parametric bivariate copulas, evaluating the $h$-functions is fast, making the simulation efficient. Increasing the dimensions of time-varying confounders, and the number of time steps does not impose significant computational challenges.
5 Demonstration
We first demonstrate our simulation method with an example replicating the setting in Bodnar et al. (2004), which examines the causal effect of iron supplement use during pregnancy on the odds of anemia at delivery.
Participants are initially randomized into treatment and control groups and are followed up in four visits leading up to delivery. Like the Iron Supplementation Study analyzed in Bodnar et al. (2004), we assume that the participants adhere to their randomly assigned treatment up to $T=3$, after which the standard clinical protocol applies.
[…]
Baseline:

### v2

@@ 3.2.2 Simulating from pair-copula constructions
For notational convenience, we introduce the h-function $h(u_{1},u_{2};\Theta)$ from Aas et al. (2009). Let $u_{1}$ and $u_{2}$ be random variables with a uniform distribution on $(0,1)$. The $h$-function is defined as
$h(u_{1},u_{2};\Theta):=F(u_{1}\,|\,u_{2})=\frac{\partial}{\partial u_{2}}C_{12}(u_{1},u_{2};\Theta)$
where the second argument in $h(\cdot;\Theta)$ is the conditional variable, and $\Theta$ represents the set of copula parameters. The $h$-function is effectively the conditional copula $C(u_{1}\,|\,u_{2};\Theta)$.
Remark 3.1.
The derivation of the analytical form of $h(\cdot)$ for common bivariate copulas is straightforward by definition for parametric families. Examples include Gaussian, Student t, Clayton, Frank, and Gumbel copulas.
The $h$-functions are foundational for simulating from $F(y\,|\,\bm{x})$. Similar to (2), by imposing an ordering $1,\ldots,d$ on the components of $\bm{x}$, we can recursively apply $h$-functions to obtain:
$F(y\,|\,\bm{x})$ $=h\{F(y\,|\,\bm{x}_{1:{d-1}}),F(x_{d}\,|\,\bm{x}_{1:{d-1}}),\Theta_{d}\}$
$=h[h\{F(y\,|\,\bm{x}_{1:{d-2}}),F(x_{d-1}\,|\,\bm{x}_{1:{d-2}};\Theta_{d-1}\},F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{d}]$
$\ldots$
$=h[\ldots h\{h(F(y),F(x_{1});\Theta_{1}),F(x_{2}\,|\,x_{1});\Theta_{2}\},\ldots,F(x_{d}\,|\,\bm{x}_{1:{d-1}});\Theta_{d}].$ (4)
Intuitively, the bivariate copulas “weave” the dependence between the outcome and each of the covariates. Thus, given a value for $\bm{X}$ we can simulate $Y$ by first drawing a uniform $U\sim{\rm U}(0,1)$, and then evaluating:
$F(Y)=h^{-1}[\ldots h^{-1}\{h^{-1}(U,\,F(X_{1});\Theta_{1}),F(X_{2}\,|\,X_{1});\Theta_{2}\},\ldots,F(X_{d}\,|\,X_{1},\ldots,X_{d-1});\Theta_{d}].$ (5)
This means that to simulate $Y$ given a $d$-dimensional vector $\bm{X}$, we recursively calculate the $d$ inverse $h$-functions, forming a pair-copula construction. This underpins our proposed simulation method.
@@ 4 Simulation Algorithm
Figure 1: An example directed acyclic graph showing temporal relationships in a longitudinal discrete setting, with baseline covariates $Z$; treatments $A_{0},\ldots,A_{K}$; time-varying covariates $L_{0},\ldots,L_{K}$ and an outcome $Y$ observed at the end of the K follow up intervals. Given the topological ordering $Z,L_{0},A_{0},L_{1},A_{1}\ldots,L_{K},A_{K},Y$, any variable may depend upon any subset of the previous variables.
We first consider a longitudinal study with $n$ individuals observed at discrete time points, $\mathcal{K}=\{0,\ldots,K\}$. At each follow-up $k$, treatment status $A_{k}$, and a set of time-varying covariates $L_{k}$ are recorded. Additionally, $Z$ represents a set of static baseline covariatesatent variables, if present, are written as $\{H_{k}\}_{k\in\mathcal{K}}$. To indicate the history of a variable up to time $k$, we use bar notation. For instance, $\bar{A}_{k}=\{A_{0},\ldots,A_{k}\}$ and $\bar{L}_{k}=\{L_{0},\ldots,L_{k}\}$, with the convention $\bar{A}_{-1}=\emptyset$ and $\bar{L}_{-1}=\emptyset$. The outcome $Y$ is observed at the final time point $K$. Figure 1 is a graphical representation of this setup, which is common in medical and epidemiological research. Examples include examining the effects of an HIV treatment on infection status assessed at the end of the study period (Robins et al., 2000), and assessing the impact of iron supplement use during pregnancy on the odds of anemia at delivery (Bodnar et al., 2004). In Section 4.4, we illustrate our method through a simulation study that replicates such settings.
@@
4.1 The fully-conditional approach
To outline our approach, we initially assume the time-varying covariate $L_{k}$ is one-dimensional and omit any latent process $\{H_{k}\}_{k\in\mathcal{K}}$; the extension to multi-dimensional time-varying covariates and latent variables is discussed in Appendix A.
The standard method for simulation of longitudinal data $(Z,\bar{L}_{K},\bar{A}_{K},Y)$ is by time step in blocks using the Rosenblatt transformation (Rosenblatt, 1952). Specifically:
@@ 4.1 The fully-conditional approach
    Baseline simulation: Simulate static baseline variables, $z$ following a specified factorization, e.g.  $F_{Z_{1}},F_{Z_{2}\mid Z_{1}},\ldots$.
@@
    Time-step simulation: At each time point $k=0,\dots,K$, simulate $l_{k}=F_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}^{-1}(w_{k,1},\,|\,z,\bar{l}_{k-1},\bar{a}_{k-1})$ and $a_{k}=F_{A_{k}\mid Z,\bar{L}_{k},\bar{A}_{k-1}}^{-1}(w_{k,2},\,|\,z,\bar{l}_{k},\bar{a}_{k-1})$, where $w_{k,1},w_{k,2}\sim\operatorname{U}(0,1)$.
@@
    Outcome simulation: Given the simulated covariates and treatments, generate the outcome $y=F_{Y\mid Z,\bar{L}_{K},\bar{A}_{K}}^{-1}(w_{y}\,|\,z,\bar{l}_{K},\bar{a}_{K})$, with $w_{Y}\sim\operatorname{U}(0,1)$.
Our proposed method diverges from this fully-conditional approach at Step 3. Instead of specifying the conditional distribution $F_{Y\mid Z\bar{L}_{K}\bar{A}_{K}}$—which inherently introduces misalignment between conditional and the desired marginal models—we introduce a margin-preserving parameterization for outcome simulation.
@@ 4.2 Frugal parameterization for outcome simulation
Our method uses the frugal parameterization to simulate outcomes, avoiding potentially intractable analytical calculations to obtain the marginal structural model. Evans and Didelez (2024) showed that the fully-conditional distribution of the outcome can be frugally parameterized as:
$p_{Y\mid\bar{L}_{K},\bar{A}_{K}}(y\,|\,\bar{l}_{k},\bar{a}_{K})=p^{*}_{Y\mid\bar{A}_{K}}(y\,|\,do(\bar{a}_{K}))\cdot c^{*}_{Y\bar{L}_{K}\mid\bar{A}_{K}}(u^{*}_{Y},u_{L_{0}},\dots,u_{L_{K}}\,|\,do(\bar{a}_{K})),$ (6)
where $u^{*}_{Y}=F_{Y\mid\bar{A}_{K}}(y\,|\,do(\bar{a}_{K}))$ and $u_{L_{k}}=F_{L_{k}\mid\bar{A}_{K}}(l_{k}\,|\,do(\bar{a}_{K}))$ for $k=0,\ldots,K$ are the marginal quantiles under interventions that set $\bar{A}_{K}=\bar{a}_{K}$.
Remark 4.1.
For brevity, we omit conditioning on baseline variables $Z$ in all quantities in (6), treating them as if no baseline variables exist. This omission will continue hereafter, though all expressions can be extended to condition on $Z$ by including it in their respective conditioning sets.
This decomposition implies that we can simulate $Y$ through the multivariate copula and its interventional margin. Applying the pair-copula construction in (2) to (6), we obtain:
$p(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=p(y\,|\,do(\bar{a}_{K}))\cdot\prod_{k=0}^{K}c^{*}\{F(y\,|\,\bar{l}_{k-1},do(\bar{a}_{K})),F(l_{k}\,|\,\bar{l}_{k-1},do(\bar{a}_{K}));\Theta_{k}\}.$ (7)
Note that we omit some subscripts on densities for conciseness where doing so does not introduce ambiguity. This decomposes the conditional density into the product of $K+1$ pair-copulas and the target marginal causal density. The pair-copula linking $Y$ and $L_{k}$ is parameterized by $\Theta_{k}$, which may depend on intervened treatments $do(\bar{A}_{k})$, past time-varying confounders $\bar{L}_{k-1}$ and baseline covariates $\bm{Z}$. Under intervention, graphically speaking, all arrows into nodes $\bar{A}_{k}$ are removed in the directed acyclic graph, rendering $L_{k}$ independent of subsequent interventions. Consequently, $F(l_{k}\,|\,\bar{l}_{k-1},do(\bar{a}_{K}))=F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$ for $k=1,\ldots,K$ and $F(l_{0}\,|\,do(\bar{a}_{K}))=F(l_{0})$. Following the simulation strategy from pair-copula constructions outlined in Section 3.2.2, the fully-conditional distribution of the outcome is alternatively parameterized as:
$F(y\,|\,\bar{l}_{K},\bar{a}_{K})$ $=h\big{[}\dots h\{h(F^{*}(y\,|\,do(\bar{a}_{K})),F(l_{0});\Theta_{0}),F(l_{1}\,|\,l_{0},a_{0});\Theta_{1}\}\dots;\Theta_{K}\big{]},$
which is centered around the causal margin $F^{*}(y\,|\,\bar{a}_{K})$.
@@
    Calculate the marginal quantile of the outcome, $u^{*}_{Y}$, as:
    $u^{*}_{Y}$ $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{Y},F(l_{K}\,|\,\bar{l}_{K-1},\bar{a}_{K-1});\Theta_{K}),$
    $\qquad\qquad\qquad F(l_{K-1}\,|\,\bar{l}_{K-2},\bar{a}_{K-2});\Theta_{K-1}\big{\}}\dots,F(l_{0});\Theta_{0}\big{]}$ (8)
    $=h^{-1}\big{[}\dots h^{-1}\big{\{}h^{-1}(w_{y},w_{K,1};\Theta_{K}),w_{K-1,1};\Theta_{K-1}\big{\}},\ldots,w_{0,1};\Theta_{0}\big{]},$
    where $w_{Y}\sim\operatorname{U}(0,1)$.
@@
    Obtain $y$ through the inverse cumulative distribution function of the marginal distribution: $y=F^{*-1}_{Y\mid\bar{A}_{K}}(u^{*}_{Y}\,|\,do(\bar{a}_{K}))$.
An intuitive way to understand the calculation in Step 1 is that the inverse $h$-functions “disentangle” the dependencies between $Y$ and each of $L_{K},L_{K-1},\dots L_{0}$ sequentially, and remove the conditioning of $Y$ on $\bar{L}_{K}$ until the marginal quantile of $Y$ is obtained. The quantiles in the inverse $h$-functions, such as $F(l_{k}\,|\,\bar{l}_{k-1},\bar{a}_{k-1})$, correspond exactly to the random standard uniform variables $w_{k,1}$ used to simulate $l_{k}$ at time step $k$. This enables direct substitution of the $l_{k}$ quantiles into (8) without additional computation.
@@ 4.3 Simulation Algorithm
In summary, to simulate the data $(Z,\bar{L}_{K},\bar{A}_{K},Y)$, the user specifies:
  • the population distribution of the static baseline covariates $Z$: $F_{Z}$, either as a joint distribution, or through a conditional factorization;
  • the development of the time-varying confounder $L_{k}$ as influenced by $\bm{Z}$ and $\bar{A}_{k-1}$: $F_{L_{k}\mid Z,\bar{L}_{k-1},\bar{A}_{k-1}}$;
  • treatment assignment based on the time-varying confounder $\bar{L}_{k}$, past treatment and baseline covariates $Z$: $F_{A_{k}\mid\bm{Z},\bar{L}_{k},\bar{A}_{k-1}}$;
  • the target causal marginal structural model: $F_{Y\mid\bar{A}_{k},Z}^{*}(\cdot\,|\,do(\bar{a}_{K}),z)$;
  • $K+1$ bivariate copulas describing dependencies between outcome $Y$ and $L_{0},\ldots,L_{K}$:
    $C_{YL_{k}\mid Z,do(\bar{A}_{k}),\bar{L}_{k-1}}(\cdot;\Theta_{k})\quad\text{for}\quad k=0,\ldots,K.$
Section 4.4 provides an example of specifying and interpreting these inputs through a simulation study that captures real-world complexities.
@@
     Sample $w_{k,1},w_{k,2}$ independent standard uniform variables
@@
Extending the simulation algorithm to include a latent process $\{H_{k}\}_{k\in\mathcal{K}}$, or multiple time-varying confounders is straightforward. The key modification lies in the outcome simulation: the marginal quantile of $Y$ is obtained by sequentially removing the conditioning on all the time-dependent confounders at each step, through inverse $h$-functions. We provide an illustrative example including a latent variable $H_{k}$ in Appendix A.
4.4 Illustrative example
We demonstrate our simulation method by replicating the setting in Bodnar et al. (2004), which examines the causal effect of iron supplement use during pregnancy on the odds of anemia at delivery. Participants are randomized into treatment and control groups and followed up over four visits leading up to delivery. Treatment is not time-fixed as participants are assumed to adhere to assigned treatment up to $k=2$, after which standard clinical protocols dictate further treatment based on hemoglobin levels.